OpenStudy (anonymous):

help please! The general formula for kinetic energy is k = 0.5mv2, where k = kinetic energy, m = mass, and v = velocity. Solve k = 0.5mv2 for v. v = m over 2k v = square root of the quantity m over 2k v = 2k over m v = square root of the quantity 2k over m

OpenStudy (anonymous):

solve for v; isolate it: \[k = \frac{ mv^2 }{ 2 }\]\[\frac{ 2k }{ m } = v^2\] so v =?

OpenStudy (anonymous):

2k over m?

OpenStudy (elise_a18):

@euler271